OpenStudy (anonymous):

(2/3)^x=9/4 I need someone to help with the steps.

OpenStudy (anonymous):

simplify your right side (2/3)^x = (3/2)^2 now, recall that to get the reciprocal, you will need to have something to the power -1 therefore, (2/3)^x = (2/3)^-2

OpenStudy (anonymous):

amendment: and thus, x = -2,